# Learning to Discover Generalized Facial Expressions

### v1

@@ 2.2 Implicit Debiasing for Theoretical Bias
Where $(\alpha-\theta)$ is an adjustment parameter, $\psi$ represents the feature generator. Considering that in Definition 1, we use the mapping function $f$ to transform the softmax probability vector into Euclidean space, we now apply $f^{-1}:\mathbb{E}\rightarrow\mathbb{P}$ to revert back to the probability space. Thus, we use cross-entropy and construct the following adversarial loss:
Estimate the F-discrepancy, corresponding to the max process of Eq. (7).
$\mathcal{L}_{ad}=\frac{1}{n}\sum_{j=1}^{n}\ell_{CE}(h_{a}(\psi(x^{u}_{j}),\hat{\mathcal{P}}^{h}(x_{j}^{u}))-\frac{\alpha}{m}\sum_{i=1}^{m}\ell_{CE}(h_{a}(\psi(x^{l}_{i}),\mathcal{F}(x_{i}^{l})),$ (8)
@@
Where $x^{u}$ and $x^{l}$ are labeled and unlabeled data, respectively, and $\hat{\mathcal{P}^{h}}$ is the pseudo-label assigned by main classification head $h$ to the unlabeled data, $\alpha$ coefficient is set to 2 according to the Tab 7. $\mathcal{L}_{ad}$ implies that approaching the bias requires $h$ and $h_{a}$ to be consistent on the labeled dataset $\mathcal{D}_{L}$, while being as inconsistent as possible on $\mathcal{D}_{U}$.
The implicit debiasing procedure has established the upper bound of bias and formulated an optimization target to minimize the maximum bias. The ’min’ component of Eq. (7) will be discussed in more detail during the explicit debiasing phase, corresponding to Eq. (12).
2.3 Explicitly Debiasing for Practical Bias
In the implicit debiasing phase, we establish a min-max optimization objective to address the Theoretical Bias of Challenge I. Additionally, for the Practical Bias of Challenge II, we aim to further enhance the performance of the feature generator and classifier to adapt to the O-FER scenario.
2.3.1 Optimization of the Feature Generator
@@
Learning Subtle Facial Features aims to acquire discriminative features from feature generator $\psi$ that enable the classifier to effectively categorize all categories. In this stage, we utilize two types of contrastive learning. For any two random augmented versions $\hat{\mathbf{x}}_{i}$ and $\tilde{\mathbf{x}}_{i}$ of an image $\mathbf{x}_{i}$ within a training batch $\mathcal{B}$ that comprises both labeled and unlabeled samples. The self-supervised contrastive loss is defined as:
$\mathcal{L}_{\text{rep}}^{u}(\hat{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{i};\psi,\tau_{u})=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}-\log\frac{\exp(\cos(\psi(\hat{\mathbf{x}}_{i}),\psi(\tilde{\mathbf{x}}_{i}))/\tau_{u})}{\sum_{\mathbf{x}_{k}\in\mathcal{B}}\exp(\cos(\psi(x_{k}),\psi(\tilde{\mathbf{x}}_{i}))/\tau_{u})},$ (9)
@@
where $\tau_{u}$ denotes a scaling parameter known as the temperature. Similar to Eq. (9), we construct supervised contrastive loss [12] $\mathcal{L}_{\text{rep}}^{s}(\hat{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{i},y;\psi,\tau_{c})$ on labeled data. These two losses are merged to shape the learning objective for the representation: $\mathcal{L}_{\text{rep}}=(1-\lambda)\mathcal{L}_{\text{rep}}^{u}+\lambda\mathcal{L}_{\text{rep}}^{s}$, with $\lambda$ serving as a tuning parameter.
Forming Discriminative Facial Expression Boundaries (Clustering). Clustering is essential for distinguishing expressions. To enhance feature discriminability, we integrate clustering with contrastive learning, which typically lacks global data structure awareness. Specifically, we connect the feature generator $\psi$ to a supervised clustering algorithm.
The clustering loss $\mathcal{L}_{\text{cluster}}$ comprises two components:
@@
The clustering loss is computed per mini-batch of labeled data $\mathcal{B}^{l}$, activated after $T_{\text{warmup}}$. Here, $\mu_{c}$ represents the class mean, $\mu_{g}$ is the global feature mean, and $n_{c}$ is the number of samples in class $c$. The set $\mathcal{C}^{l}$ includes all unique class labels in the batch, while $\mathcal{B}^{l}_{c}$ denotes the samples belonging to class $c$. The overall clustering loss is defined as $\mathcal{L}_{\text{cluster}}=\mathcal{L}_{\text{WB}}+\beta\cdot\mathcal{L}_{\text{MM}}$. The term $\mathcal{L}_{\text{WB}}$ aims to minimize intra-class distances and maximize inter-class distances, while $\mathcal{L}_{\text{MM}}$, as a regularization term, promotes compact clustering by reducing the gap between maximum and minimum distances.
2.3.2 Fine-tuning of Main Classification Head
@@
Parametric classifier learning has become a effective paradigm. We use it to augment main head $h$. We follow the SimGCD [10] approach to assign labels to instance inputs. Specifically, the number of categories $K=\mathcal{M}+\mathcal{N}$ is given. A set of parametric prototypes for each category $\mathcal{T}=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{K}\}$ is randomly initialized at the beginning. During training, the soft label $\hat{\mathbf{p}}_{i}^{k}$ for each augmented view $\mathbf{x}_{i}$ is calculated using a softmax function on the cosine similarity between the hidden feature and the prototypes:
$\hat{\mathbf{p}}_{i}^{k}=\frac{\exp\left(\frac{1}{\tau_{s}}(h(\psi(\mathbf{x}_{i}))/{\|h(\psi(\mathbf{x}_{i}))\|_{2}})^{\top}({\mathbf{t}_{k}}/{\|\mathbf{t}_{k}\|_{2}})\right)}{\sum_{j}\exp{\left(\frac{1}{\tau_{s}}(h(\psi(\mathbf{x}_{i}))/\|h(\psi(\mathbf{x}_{i}))\|_{2})^{\top}(\mathbf{t}_{j}/\|\mathbf{t}_{j}\|_{2})\right)}},$ (11)
@@
Similarly, we can obtain the soft label $\tilde{\mathbf{p}}_{i}$ of the view $\tilde{\mathbf{x}}_{i}$. The supervised and unsupervised losses of the classifier are formulated by:
$\mathcal{L}_{\text{cls}}^{s}$ $=\frac{1}{|\mathcal{B}^{l}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}^{l}}\ell_{CE}(\mathbf{y}({x_{i}}),\hat{\mathbf{p}}_{i}),\,\mathcal{L}_{\text{cls}}^{u}=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}\ell_{CE}(\tilde{\mathbf{p}}_{i},\hat{\mathbf{p}}_{i})-\epsilon H(\overline{\mathbf{p}}),$ (12)
@@
where $\mathcal{B}^{l}$ is the mini-batch of labeled training data, $\mathbf{y}({x_{i}})$ is the ground truth label for the labeled data point $\mathbf{x}_{i}$, $\ell_{CE}$ is the cross-entropy loss, and $H(\overline{\mathbf{p}})=-\sum\overline{\mathbf{p}}\log\overline{\mathbf{p}}$ regularizes the mean prediction $\overline{\mathbf{p}}=\frac{1}{2|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}(\hat{\mathbf{p}}_{i}+\tilde{\mathbf{p}}_{i})$ in a mini-batch. Then the objective of parametric classifier learning is $\mathcal{L}_{\text{cls}}=(1-\lambda)\mathcal{L}_{\text{cls}}^{u}+\lambda\mathcal{L}_{\text{cls}}^{s}$.
Elimination of imbalance bias. In FER scenarios, sample imbalance often biases models toward majority classes, reducing accuracy for minority classes. This issue is further complicated when the minority class distribution is unknown. To address this, we introduce a dynamic reconciliation process that adaptively adjusts training, improving model generalization. We achieve this by utilizing the following loss function:
$\mathcal{L}_{\text{bal}}(\mathbf{x},y;\psi,h)=-\sum_{x_{i}\in\mathcal{B}}\sum_{c\in\mathcal{C}}\delta(y_{i},c)\left[(1-e_{i})+\frac{e_{i}}{a_{c}+\epsilon}\right]\log\left(\frac{\exp(h(\psi(\mathbf{x}_{i}))_{c})}{\sum_{k\in\mathcal{C}}\exp(h(\psi(\mathbf{x}_{i}))_{k})}\right),$ (13)
Where $a_{c}$ is the adaptive weight for class $c$, calculated as the ratio of $\varsigma_{c}$ (the number of correct predictions for class $c$) to $\eta_{c}$ (the total number of predictions made for class $c$), i.e., $a_{c}=\frac{\varsigma_{c}}{\eta_{c}}$. The time-varying weight $e_{i}$ decreases as training progresses, given by $e_{i}=0.1\cdot\left(1-\frac{t}{T}\right)$, where $t$ is the current training round and $T$ is the total rounds. This method dynamically tracks poorly predicted categories, eliminating the need for prior knowledge about the number of classes.
Overall loss function: Through the aforementioned process, we have defined the maximum bias and enhanced the model’s robustness in the presence of the maximum bias to eliminate its impact. The overall loss function is as follows:
$\mathcal{L}_{total}=\mathcal{L}_{rep}+\mathcal{L}_{cls}-\lambda_{a}\mathcal{L}_{ad}+\lambda_{b}\mathcal{L}_{bal}+\lambda_{c}\mathcal{L}_{cluster}$ (14)
@@
Where $\lambda_{a}$, $\lambda_{b}$ and $\lambda_{c}$ are three balancing parameters, we set them to 0.2, 0.3, and 0.2, respectively, based on the hyperparameter analysis in appendix B.3. $\mathcal{L}_{cluster}$ is only introduced after the warmup epoch $T_{warmup}$, when the model has preliminary clustering capabilities.
Table 1: Category discovery accuracy (ACC) on RAF-DB containing only known expressions. We test Ada-CM under GCD setting. $\triangle$ represents the improvement accuracy.

### v2

@@ 4.1 Implicit Debiasing
where $(\alpha-\theta)$ is an adjustment parameter, $\psi$ represents the feature exactor. Considering that in Definition 1, we use the mapping function $f$ to transform the softmax probability vector into Euclidean space, we now apply $f^{-1}:\mathbb{E}\rightarrow\mathbb{P}$ to revert back to the probability space. Thus, we use cross-entropy and construct the following adversarial loss:
$\mathcal{L}_{ad}=\frac{\alpha}{m}\sum_{i=1}^{m}\ell_{CE}$ $\left(h_{a}(\psi(x^{l}_{i})),h(\psi(x^{l}_{i}))\right)$ (8)
$-\frac{1}{n}\sum_{j=1}^{n}$ $\ell_{CE}\left(h_{a}(\psi(x^{u}_{j})),\hat{\mathcal{P}}^{h}(x_{j}^{u})\right),$
@@
where $x^{u}$ and $x^{l}$ are labeled and unlabeled data, respectively, and $\hat{\mathcal{P}^{h}}$ is the pseudo-label assigned by main classification head $h$ to the unlabeled data, $\alpha$ coefficient is set to 2 according to the hyperparameter analysis in the appendix.
4.2 Explicitly Debiasing
Explicit bias arises from overlapping characteristics between known and unknown categories, leading to ambiguous boundaries. To address this, we propose a hierarchical category-discrimination refinement strategy through sample-level, triplet-level, and distribution-level optimizations, complemented by a parametric classifier to enhance decision-making in the G-FACE task.
Sample-level: Confusing Sample Mining. To address ambiguity in overlapping characteristics, we propose a dynamic adjustment mechanism applied to reach labeled data $\mathcal{B}^{l}$. This strategy focuses on mining confusing samples and enhancing clarity in known categories while preserving the model’s ability to discover new categories in unlabeled data. Without requiring prior knowledge of class distribution, the loss function is defined as:
$\mathcal{L}_{\text{bal}}(\mathbf{x},y;\psi,h)=\frac{1}{|\mathcal{B}^{l}|}\sum_{x_{i}\in\mathcal{B}^{l}}\left[(1-e_{i})+\frac{e_{i}}{a_{y_{i}}+\epsilon}\right]\ell_{CE}(h(\psi(\mathbf{x}_{i})),y_{i}),$ (9)
where $a_{y_{i}}$ is the adaptive weight for the true class $y_{i}$, calculated as the ratio of $\varsigma_{y_{i}}$ (the number of correct predictions for class $y_{i}$) to $\eta_{y_{i}}$ (the total number of predictions made for class $y_{i}$), i.e., $a_{y_{i}}=\frac{\varsigma_{y_{i}}}{\eta_{y_{i}}}$. The time-varying weight $e_{i}$ decreases as training progresses, given by $e_{i}=0.1\cdot\left(1-\frac{t}{T}\right)$, where $t$ is the current epoch and $T$ is the epoch. This mechanism dynamically emphasizes underperforming classes, improving the clarity of decision boundaries. Furthermore, for the dynamically adjusted $e_{i}$, we conduct experimental validation in the appendix.
@@ 4.2 Explicitly Debiasing
Triplet-level: Learning Subtle Facial Expression Features. This level focuses on acquiring discriminative features from the feature extractor $\psi$ based on a triplet of samples, enabling the classifier to effectively distinguish between all categories. To this end, we utilize two types of contrastive learning (CL), including self-supervised CL and supervised CL. Using any two random augmented versions $\hat{\mathbf{x}}_{i}$ and $\tilde{\mathbf{x}}_{i}$ of a face image $\mathbf{x}_{i}$ within a training batch $\mathcal{B}$ that comprises both labeled and unlabeled samples, the self-supervised CL loss is defined as:
$\mathcal{L}_{\text{rep}}^{u}(\hat{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{i};\psi,\tau_{u})$ $=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}-\log\frac{\exp(\hat{\mathbf{z}}_{i}^{\top}\tilde{\mathbf{z}}_{i}/\tau_{u})}{\sum_{\mathbf{x}_{k}\in\mathcal{B}}\exp(\hat{\mathbf{z}}_{k}^{\top}\tilde{\mathbf{z}}_{i}/\tau_{u})},$ (10)
@@
where $\tau_{u}$ denotes a scaling parameter known as the temperature, $\hat{z}_{i}$ is the $\ell_{2}$-normalised vector of $\psi(\hat{\mathbf{x}}_{i})$. Similarly, we construct the supervised CL loss [15] as $\mathcal{L}_{\text{rep}}^{s}(\hat{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{i},y;\psi,\tau_{c})$ on labeled data. These two losses are merged to learn the subtle facial expression representation as: $\mathcal{L}_{\text{rep}}=(1-\lambda)\mathcal{L}_{\text{rep}}^{u}+\lambda\mathcal{L}_{\text{rep}}^{s}$, with $\lambda$ as a hyper-parameter based on previous GCD study [35].
Distribution-level: Enhancing Cluster Discriminability. To improve the discriminability of expression boundaries, we combine clustering with contrastive learning, which alone often lacks awareness of the global data structure. Specifically, we enhance the feature exactor $\psi$ by incorporating a supervised clustering loss. The clustering loss $\mathcal{L}_{\text{cluster}}$ comprises two components as $\mathcal{L}_{\text{cluster}}=\mathcal{L}_{\text{WB}}+\beta\cdot\mathcal{L}_{\text{MM}}$, where $\beta$ is the weight of regular term and set as 0.2 in this study. The term $\mathcal{L}_{\text{WB}}$ aims to minimize intra-class distances and maximize inter-class distances, while $\mathcal{L}_{\text{MM}}$, as a regularization term, promotes compact clustering by reducing the gap between maximum and minimum distances. Formally, the $\mathcal{L}_{\text{WB}}$ and $\mathcal{L}_{\text{MM}}$ are written as:
@@
where $\mu_{c}$ represents the class mean, $\mu_{g}$ is the global feature mean, and $n_{c}$ is the number of samples in class $c$. The set $\mathcal{C}^{l}$ includes all unique class labels in the batch, while $\mathcal{B}^{l}_{c}$ denotes the samples belonging to class $c$. The total $\mathcal{L}_{\text{cluster}}$ is computed per mini-batch of labeled data $\mathcal{B}^{l}$, activated after $T_{\text{warmup}}$.
@@
Parametric Classifier Learning. The parametric classifier is used to augment the main expression classification head $h$. To enhance the main expression classification head $h$, we employ a parametric classifier following the approach of SimGCD [39]. Specifically, the number of categories $K=\mathcal{M}+\mathcal{N}$ is given. We start by randomly initializing a set of prototypes, one for each category: $\mathcal{T}=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{K}\}$ is randomly initialized at the beginning. During training, the soft label $\hat{\mathbf{p}}_{i}^{k}$ for each augmented view $\mathbf{x}_{i}$ is calculated using a softmax function on the cosine similarity between the hidden feature and the prototypes:
$\small\hat{\mathbf{p}}_{i}^{k}=\frac{\exp\left(\frac{1}{\tau_{s}}(h(\psi(\mathbf{x}_{i}))/{\|h(\psi(\mathbf{x}_{i}))\|_{2}})^{\top}({\mathbf{t}_{k}}/{\|\mathbf{t}_{k}\|_{2}})\right)}{\sum_{j}\exp{\left(\frac{1}{\tau_{s}}(h(\psi(\mathbf{x}_{i}))/\|h(\psi(\mathbf{x}_{i}))\|_{2})^{\top}(\mathbf{t}_{j}/\|\mathbf{t}_{j}\|_{2})\right)}}.$ (12)
@@
Similarly, we can obtain the soft label $\tilde{\mathbf{p}}_{i}$ of the view $\tilde{\mathbf{x}}_{i}$. Then the objective of parametric classifier learning is the sum of the supervised loss and unsupervised loss as $\mathcal{L}_{\text{cls}}=(1-\lambda)\mathcal{L}_{\text{cls}}^{u}+\lambda\mathcal{L}_{\text{cls}}^{s}$. The supervised and unsupervised losses are formulated by:
$\mathcal{L}_{\text{cls}}^{s}=\frac{1}{|\mathcal{B}^{l}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}^{l}}\ell_{CE}(\mathbf{y}({x_{i}}),\hat{\mathbf{p}}_{i}),\,$ (13)
$\mathcal{L}_{\text{cls}}^{u}=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}\ell_{CE}(\tilde{\mathbf{p}}_{i},\hat{\mathbf{p}}_{i})-\epsilon H(\overline{\mathbf{p}}),$
@@
where $\mathcal{B}^{l}$ is the mini-batch of labeled training data, $\mathbf{y}({x_{i}})$ is the ground truth label for the labeled data point $\mathbf{x}_{i}$, $\ell_{CE}$ is the cross-entropy loss, and $H(\overline{\mathbf{p}})=-\sum\overline{\mathbf{p}}\log\overline{\mathbf{p}}$ regularizes the mean prediction $\overline{\mathbf{p}}=\frac{1}{2|\mathcal{B}|}\sum_{\mathbf{x}_{i}\in\mathcal{B}}(\hat{\mathbf{p}}_{i}+\tilde{\mathbf{p}}_{i})$ in a mini-batch.
4.3 Overall Loss Function:
The overall loss function in our DIG-FACE is as follows:
$\mathcal{L}_{total}=\mathcal{L}_{rep}+\mathcal{L}_{cls}+\lambda_{a}\mathcal{L}_{ad}+\lambda_{b}\mathcal{L}_{bal}+\lambda_{c}\mathcal{L}_{cluster},$ (14)
@@ 4.3 Overall Loss Function:
where $\lambda_{a}$, $\lambda_{b}$ and $\lambda_{c}$ are balancing parameters for the contributions of two debasing processes during training, we set them to 0.2, 0.3, and 0.2, respectively, based on the hyperparameter analysis in our supplementary appendix. $\mathcal{L}_{cluster}$ is only introduced after the warmup epoch $T_{\text{warmup}}$, when the model has preliminary clustering capabilities.
Table 2: Category discovery accuracy (ACC) on RAF-DB dataset. Above the table are the four selected known categories.